n66 13 2

Evaluate the expression.
Answer #1

n C x = n! / x!(n - x)!

n! = n(n - 1)(n - 2) .... and so on .

6 C 1 = 6

5 C 3 = 10

2 C 2 = 1

Expression = 6 * 10 * 1 = 60

media%2F3e4%2F3e484a73-9e66-476b-814e-71 = 60